Charlie’s Furniture Store has been in business for several years. The firm's owners have described the store as a "high-price, high-service" operation that provides lots of assistance to its customers. Margin has averaged a relatively high 32% per year for several years, but turnover has been a relatively low 0.4 based on average total assets of $1,200,000. A discount furniture Store is about to open in the area served by Charlie's, and management is considering lowering prices to compete effectively.

Required:

  1. Calculate current sales and ROI for Charlie’s Furniture Store.
  2. Assuming that the new strategy would reduce margin to 20%, and assuming that average total assets would stay the same, calculate the sales that would be required to have the same ROI as Charlie’s currently earns.
  3. Suppose you presented the results of your analysis in parts a and b of this problem to Charlie, and he replied, “What are you telling me? If I reduce my prices as planned, then I have to increase my sales volume by 50% to earn the same return?” Given the results of your analysis, what is the actual amount of increase in sales required?
  4. Now suppose Charlie says, “You know, I’m not convinced that lowering prices is my only option in staying competitive. What if I were to increase my marketing effort? I’m thinking about kicking off a new advertising campaign after conducting more extensive market research to better identify who my target customer groups are.” In general, explain to Charlie what the likely impact of a successful strategy of this nature would be on margin, turnover, and ROI.
  5. What are the other alternative strategy that might help Charlie maintain the competitiveness of his business.
Answer #1
  1. Current Sales = $1200000*0.4 = $480000

ROI = $480000*32%/$1200000 = 12.8%

  1. Sales if Margin is reduced by 20% and ROI remains same

Sales = 12.8%*$1200000/20% = $768000

  1. The Actual Amount of Increase in Sales would be = $768000

The increase in sales volume to earn same return should be 60%

  1. In this scenario Margin would be decrease and advertisement cost would increase.

Turnover would increase and ROI would be decrease as Margin would be reduced.

  1. Increase in selling price, Labour saving strategies, Reduction in inventory carrying costs.